# Cycles in spherical Deligne complexes and application to $K(\pi,1)$-conjecture for Artin groups

### v1

@@ 9 $K(\pi,1)$-conjecture for some 3D Artin groups
Let $A_{S}$ be an Artin group with $|S|=4$. Suppose its presentation graph $\Gamma$ is a complete graph. If there exists $S^{\prime}\subset S$ with $|S^{\prime}|=3$ such that $A_{S^{\prime}}$ is not spherical, then the Artin complex for $A_{S}$ is contractible.
@@
Let $S=\{a,b,c,d\}$ and $S^{\prime}=\{a,b,c\}$. Let $\Delta_{S}$ (resp. $\Delta_{S^{\prime}}$) be the Artin complex for $A_{S}$ (resp. $A_{S^{\prime}}$). By Lemma 3.3, the link of each vertex of type $\hat{d}$ in $\Delta_{S}$ is isomorphic to $\Delta_{S^{\prime}}$. As $A_{S^{\prime}}$ is not spherical, it is a 2-dimensional Artin group. It is known the modified Deligne complex, defined in [CD95a], for any 2-dimensional Artin group is contractible [CD95a]; and the modified Deligne complex isomorphic to the barycentric subdivision of $\Delta_{S^{\prime}}$. Thus the link of each vertex of type $\hat{d}$ is contractible. Then $\Delta_{S}$ is homotopic equivalent to the relative Artin complex $\Delta^{\prime}=\Delta_{S,S^{\prime}}$ by Lemma 3.5.
@@
Now we look at the case $A_{bcd}$ is not spherical. Then $\operatorname{lk}(w,\Delta_{S})$ quotiented by the action of $A_{bcd}$ is a triangle $T^{\prime}$, whose vertices are denoted by $\hat{b},\hat{c},\hat{d}$. As $\frac{1}{m_{bc}}+\frac{1}{m_{cd}}+\frac{1}{m_{bd}}\leq 1$, we can realize $T^{\prime}$ as a geodesic triangle in $\mathbb{E}^{2}$ or $\mathbb{H}^{2}$ such that the angle at $\hat{b}$ is $\frac{\pi}{m_{cd}}$, the angle at $\hat{c}$ is $\frac{\pi}{m_{bd}}$ and the angle at $\hat{d}$ is $\frac{\pi}{m_{bc}}$. This induces a metric on $\operatorname{lk}(w,\Delta_{S})$ which is known to be CAT$(0)$ (cf. [CD95a]). As $\operatorname{lk}(w,\Delta^{\prime})$ is the full subcomplex of $\operatorname{lk}(w,\Delta)$ spanned by vertices of type $\hat{b}$ and $\hat{c}$, we know $\operatorname{lk}(w,\Delta^{\prime})$ has girth $2m_{bc}$ by [Hua23, Lemma 9.8 (1)]. Thus the claim is proved. ∎
@@
Suppose $A_{S}$ is an Artin group with $|S|\leq 4$. Assume that its Dynkin diagram is not a $(3,5,3)$-linear diagram. Then $A_{S}$ satisfies the $K(\pi,1)$-conjecture.
@@
We assume $A_{S}$ is irreducible. By [ES10], it suffices to consider the case when the presentation graph of $A_{S}$ is complete. We also assume $A_{S}$ is not spherical, otherwise the result follows from [Del72]. When $|S|\leq 3$, $A_{S}$ is 2-dimensional and the theorem follows from [CD95b]. Now we assume $|S|=4$. If there exists a subset $S^{\prime}\subset S$ such that $|S^{\prime}|=3$ and $A_{S^{\prime}}$ is not spherical, then the theorem follows from Proposition 9.1. It remains to consider the case when $|S|=4$ and $A_{S^{\prime}}$ is spherical whenever $S^{\prime}$ is a 3-element subset of $S$. By a result of Lanner [Lan50a], $A_{S}$ is either an affine Artin group, or a 3-dimensional hyperbolic cocompact tetrahedron group, i.e. its associated Coxeter group acts on $\mathbb{H}^{3}$ properly and cocompactly by isometries such that its fundamental domain is a tetrahedron. The $K(\pi,1)$-conjecture for affine Artin groups are proved in [PS21]. For the 3-dimensional hyperbolic tetrahedron groups, there are only nine of them, five has Dynkin diagram being a cycle, which is treated in [garside, Hua23]. Then remaining four groups have Dynkin diagrams as in Figure 15. By Theorem 3.1, it suffices to show the associated Artin complex is contractible.
@@
By Theorem 3.1, it suffices to show the Artin complexes associated with these Dynkin diagrams are contractible. First we look at the case of $(5,3,4)$ and $(5,3,5)$ linear diagrams. Assume the vertices in the Dynkin diagram are ordered from left to right, and we endow the vertex set of the associated Artin complex $\Delta$ with the induced order. Then the contractibility of $\Delta$ follows from Theorem 3.27, Theorem 7.1, Theorem 3.22, Theorem 3.17 and Lemma 3.23. In the case when $\Lambda$ is a tripod, we deduce the contractibility of the Artin complex from Proposition 3.35, Theorem 7.1, Theorem 5.6, and Theorem 3.17. ∎
@@
Let $W_{\Gamma}$ be a reflection group acting properly on $\mathbb{H}^{3}$ with finite volume fundamental domain, with its presentation graph being $\Gamma$. Assume the Dynkin diagram of $W_{\Gamma}$ is not the linear graph with consecutive edges labeled by $(3,5,3)$. Then the $K(\pi,1)$-conjecture holds true for the associated Artin group $A_{\Gamma}$.

### v3

@@ 9.2 Case $\widetilde{D}_{n}$
9.3 Downward flagness with respect to different subdivisions
Throughout this subsection, let $\Lambda,\Lambda^{\prime},\{\Lambda_{i}\}_{i=1}^{3},\{\Lambda^{\prime}_{i}\}_{i=1}^{3}$, $\Delta$ and $\Delta^{\prime}$ be as in Proposition 8.3. Our main goal in this subsection is to prove:
Proposition 9.3.
Lemma 9.4.
Let $x_{1}x_{2}x_{3}x_{4}$ be an embedded four cycle in $\Delta$ such that $\operatorname{Type}(x_{1})=\operatorname{Type}(x_{3})=\hat{b}_{i}$, $\operatorname{Type}(x_{2})=\hat{b}_{j}$, $\operatorname{Type}(x_{4})=\hat{b}_{k}$ with $\{i,j,k\}$ being pairwise distinct. Then $x_{2}\sim_{\Delta}x_{4}$.
Proof.
By Proposition 9.1 and [Hua24a, Proposition 6.15], the complex $\Delta_{\Lambda,\Lambda^{\prime}}$ satisfies the labeled four wheel condition in the sense of [Hua24a, Definition 6.12]. Thus if $x_{2}$ and $x_{4}$ are not adjacent, then there is a vertex $y\in\Delta$ of type $\hat{a}$ such that $y$ is adjacent to each of $x_{i}$. As $\Lambda^{\prime}$ is admissible in $\Lambda$, Corollary 2.3 implies $x_{2}\sim_{\Delta}x_{4}$, contradiction. ∎
Lemma 9.5.
Let $x_{1}y_{1}x_{2}y^{\prime}_{2}x_{3}y^{\prime}_{3}x_{4}y_{4}$ be consecutive vertices of a 8-cycle in $\Delta^{\prime}$ such that $\operatorname{Type}(x_{1})=\hat{b}_{3}$, $t(x_{3})=1$; $\operatorname{Type}(x_{i})=\hat{b}_{2}$ for $i=2,4$; $t(y_{i})=2$ for $i=1,4$ and $t(y^{\prime}_{i})=4$ for $i=2,3$. Then at least one of the following holds true:
  1. 1.
    there exists a vertex $z\in\Delta$ with $t(z)=4$ such that $z$ is adjacent in $\Delta$ to both $x_{2}$ and $x_{4}$;
  2. 2.
    $x_{1}$ is adjacent in $\Delta$ to both $y^{\prime}_{2}$ and $y^{\prime}_{3}$;
  3. 3.
    there exists a vertex $z\in\Delta$ of type $\hat{b}_{3}$ such that it is adjacent in $\Delta$ to each of $\{x_{2},y^{\prime}_{2},y^{\prime}_{3},x_{4}\}$.
Proof.
We will assume $x_{2}\neq x_{3}$ and $x_{4}\neq x_{3}$, otherwise we are in Case 1. Assume $x_{1}\neq x_{3}$, otherwise we are in Case 2. Assume $y^{\prime}_{2}\neq y^{\prime}_{3}$, otherwise we are in case 1. Repeating the argument in the proof of Proposition 9.1 (2) (see Figure 12, though we caution the reader that now we are not assuming $x_{3}$ is adjacent in $\Delta$ to $x_{2}$ and $x_{4}$), we can find $y^{\prime}_{1}$ with $t(y^{\prime}_{1})=4$ such that $y^{\prime}_{1}$ is adjacent in $\Delta$ to both $x_{1}$ and $x_{2}$, $y^{\prime}_{4}$ with $t(y^{\prime}_{4})=4$ such that $y^{\prime}_{4}$ is adjacent in $\Delta$ to both $x_{1}$ and $x_{4}$, and $z$ with $t(z)=1$ such that $z$ is adjacent in $\Delta$ to each of $\{y^{\prime}_{1},y^{\prime}_{2},y^{\prime}_{3},y^{\prime}_{4}\}$. We assume $y^{\prime}_{1}\neq y^{\prime}_{4}$, otherwise we are in case 1.
If $y^{\prime}_{1}=y^{\prime}_{2}$, then $\{x_{1},x_{3},x_{4}\}$ pairwise have a upper bound in $(V\Delta^{\prime},<)$. Thus they have a common upper bound by Lemma 5.1, denoted by $w$. We can assume $t(w)=4$. If $w=y^{\prime}_{2}$, then we are in case 1. Now suppose $w\neq y^{\prime}_{2}$. If $\operatorname{Type}(x_{3})=\hat{b}_{2}$, then by applying Lemma 9.4 to $x_{1}y^{\prime}_{2}x_{3}w$, we know $x_{1}\sim_{\Delta}x_{3}$. Applying Lemma 9.4 to $x_{1}x_{3}y^{\prime}_{3}x_{4}$ (which is embedded as $x_{3}\neq x_{4}$), we know $x_{1}\sim_{\Delta}y^{\prime}_{3}$, hence we are in case 2. Suppose $\operatorname{Type}(x_{3})=\hat{b}_{3}$. By considering the $4$-cycle $x_{1}y^{\prime}_{2}x_{3}w$ (which is embedded as $x_{1}\neq x_{3}$) and apply Proposition 9.1, we know there exists a vertex $z^{\prime}$ of type $\hat{a}$ such that $z^{\prime}$ is adjacent in $\Delta$ to each vertex of this 4-cycle. If $w=y^{\prime}_{3}$, then we are in Case 2. If $w\neq y^{\prime}_{3}$, then by applying Lemma 9.4 to $wx_{3}y^{\prime}_{3}x_{4}$, we know $x_{3}\sim_{\Delta}x_{4}$. By considering the 4-cycle $x_{1}z^{\prime}x_{3}x_{4}$ and Proposition 9.1 (1), we know $z^{\prime}\sim_{\Delta}x_{4}$. Thus $y^{\prime}_{2}\sim_{\Delta}x_{4}$ and we are in case 1. A similar analysis can be applied if $y^{\prime}_{3}=y^{\prime}_{4}$. Thus we assume $y^{\prime}_{i}\neq y^{\prime}_{i+1}$ for all $i$ from now on.
If $\operatorname{Type}(z)=\hat{b}_{2}$, then by Lemma 9.4, $x_{1}\sim_{\Delta}z$. Assume $z\neq x_{2}$ and $z\neq x_{4}$, otherwise we are in case 1. By applying Lemma 9.4 to the embedded 4-cycles $x_{1}zy^{\prime}_{2}x_{2}$ and $x_{1}zy^{\prime}_{3}x_{4}$ in $\Delta$, we know that $x_{1}\sim_{\Delta}y^{\prime}_{i}$ for $i=2,3$. Suppose $\operatorname{Type}(z)=\hat{b}_{3}$. Applying Lemma 9.4 to the embedded 4-cycles $zy^{\prime}_{2}x_{2}y^{\prime}_{1}$ and $zy^{\prime}_{2}x_{4}y^{\prime}_{4}$, we know $z\sim_{\Delta}x_{i}$ for $i=2,4$, hence we are in case 3. ∎
Proof of Proposition 9.3.
The case of $(b_{2},b_{3})$-subdivision follows from Lemma 5.1. The cases of $(b_{1},b_{2})$-subdivision and $(b_{1},b_{3})$-subdivision are symmetric, so we only treat the $(b_{1},b_{3})$-subdivision $\Delta^{\prime\prime}$ of $\Delta$. Let $t^{\prime\prime}$ be the type function on $V\Delta^{\prime\prime}$. By Lemma 2.22, $(V\Delta^{\prime\prime},<)$ is a poset. The bowtie free property of $(V\Delta^{\prime\prime},<)$ follows from that $\Delta_{\Lambda,\Lambda^{\prime}}$ satisfies the labeled four wheel condition (see Lemma 9.4).
Take three different vertices $\{x_{1},x_{3},x_{5}\}$ in $\Delta^{\prime\prime}$ such that they pairwise have a lower bound in $(V\Delta^{\prime\prime},<)$. We need to show they have a common lower bound. We will only treat the case when $t^{\prime\prime}(x_{i})=4$ for $i=1,3,5$; as the other cases are similar and much simpler. Let $x_{i+1}$ be a lower bound for $\{x_{i},x_{i+2}\}$, with $i=1,3,5\in\mathbb{Z}/6\mathbb{Z}$. We can assume $t^{\prime\prime}(x_{i})=1$ for $i=2,4,6$. Then the vertices $\{x_{i}\}_{i=1}^{6}$ form a 6-cycle in $\omega$. We can assume this 6-cycle is embedded, otherwise we are already done. Note that $\omega$ is also a 6-cycle in $\Delta$. Let $m_{i}\in\Delta^{\prime}$ be the midpoint of $\overline{x_{i}x_{i+1}}$. Note that $\operatorname{Type}(x_{i})=\hat{b}_{2}$ for $i$ odd. And $\operatorname{Type}(x_{i})=\hat{b}_{1}$ or $\hat{b}_{3}$ for $i$ even.
Case 1: all of $\{x_{2},x_{4},x_{6}\}$ has type $\hat{b}_{1}$. Then $\{x_{1},x_{3},x_{5}\}$ are pairwisely upper bounded in the poset $(V\Delta^{\prime},<)$. By Lemma 5.1 and Proposition 8.3, there is a vertex $z\in V\Delta^{\prime}$ which is a common upper bound for $\{x_{1},x_{3},x_{5}\}$ in $(V\Delta^{\prime},<)$. We can assume $t(z)=4$. Then $t^{\prime\prime}(z)=1$ and $z$ is a common lower bound for $\{x_{1},x_{3},x_{5}\}$ in $(V\Delta^{\prime\prime},<)$.
Case 2: exactly two of $\{x_{2},x_{4},x_{6}\}$, say $x_{2}$ and $x_{6}$, have type $\hat{b}_{1}$. Then Lemma 9.5 applies to the $8$-cycle $x_{5}x_{6}x_{1}x_{2}x_{3}m_{3}x_{4}m_{4}$ in $\Delta^{\prime}$. If we are in case 1 of Lemma 9.5, then we reduce to the previous paragraph. If we are in case 2 of Lemma 9.5, then $x_{4}$ is adjacent in $\Delta$ to $x_{6}$ and $x_{2}$. Applying Lemma 9.4 to the $4$-cycle $x_{4}x_{6}x_{1}x_{2}$ in $\Delta$, we know $x_{4}\sim_{\Delta}x_{1}$. Thus $x_{4}$ is a common lower bound of $\{x_{1},x_{3},x_{5}\}$ in $(V\Delta^{\prime\prime},<)$. If we are in case 3 of Lemma 9.5, then let $z$ be the vertex of type $\hat{b}_{3}$ such that $z\sim_{\Delta}\{x_{5},x_{6},x_{2},x_{3}\}$. Applying Lemma 9.4 to the $4$-cycle $zx_{6}x_{1}x_{2}$ in $\Delta$, we know $z\sim_{\Delta}x_{1}$. Thus $z$ is a common lower bound of $\{x_{1},x_{3},x_{5}\}$ in $(V\Delta^{\prime\prime},<)$.
Before we discuss the remaining cases, we need an extra observation. Let $Y$ be the graph as in the proof of Proposition 9.1. As consecutive vertices of $\omega$ has distance $\leq 2$ in $\Delta^{\prime}$, hence the same holds in $Y$. Thus $\{B_{Y}(x_{i},2)\}_{i=1,3,5}$ pairwise intersects. Let $z\in Y$ be a vertex in their common intersection. Then for $i=1,3,5$, either $z\sim_{Y}x_{i}$ or there is $z_{i}\in Y$ with $z_{i}\notin\{z,x_{i}\}$ such that $z_{i}\sim_{Y}x_{i}$ and $z_{i}\sim_{Y}z$. If $t(z)=4$, then the definition of $Y$ implies that for $i=1,3,5$, either $z\sim_{\Delta^{\prime}}x_{i}$, or $z_{i}\sim_{\Delta^{\prime}}x_{i}$ and $z_{i}\sim_{\Delta^{\prime}}z$. In the latter case, $t(x_{i})<t(z_{i})<t(z)$, hence $z\sim_{\Delta^{\prime}}x_{i}$ as well. Thus $z$ is a common lower bound for $\{x_{1},x_{3},x_{5}\}$ in $(V\Delta^{\prime\prime},<)$. Now we assume $t(z)<4$. Then by an argument in the proof of Proposition 9.1 (2), for each $i=1,3,5$, we can assume $t(z_{i})=4$ such that $z_{i}\sim_{\Delta^{\prime}}x_{i}$ and $z_{i}\sim_{\Delta^{\prime}}z$; moreover, up to replacing $z$, we can assume $t(z)=1$. Then $\operatorname{Type}(z)=\hat{b}_{2}$ or $\hat{b}_{3}$.
Case 3: exactly one of $\{x_{2},x_{4},x_{6}\}$, say $x_{6}$, has type $\hat{b}_{1}$. First we look at the situation that there exists a vertex $x^{\prime}_{6}$ of type $\hat{b}_{1}$ such that $x^{\prime}_{6}$ is adjacent in $\Delta$ to $\{x_{5},z,x_{1}\}$. See Figure 13 (I). Now we apply Lemma 9.5 to the 8-cycles:
$\omega=x^{\prime}_{6}zz_{3}x_{3}m_{2}x_{2}m_{1}x_{1},\ \omega^{\prime}=x^{\prime}_{6}zz_{3}x_{3}m_{3}x_{4}m_{4}x_{5}.$
If at least one of $\omega$ and $\omega^{\prime}$ are in Lemma 9.5 (1), then we are reduced to the previous cases. If none of the $\omega$ and $\omega^{\prime}$ are in Lemma 9.5, then up to possibly replacing $x_{2}$ and $x_{4}$ by different vertices of type $\hat{b}_{3}$, we can assume $x_{2}\sim_{\Delta}\{x_{1},x^{\prime}_{6},z_{3},x_{3}\}$ and $x_{4}\sim_{\Delta}\{x_{3},x^{\prime}_{6},z_{3},x_{5}\}$. As $x_{2}\neq x_{4}$ (otherwise we are done), by applying Lemma 9.4 to the 4-cycle $x^{\prime}_{6}x_{2}x_{3}x_{4}$ in $\Delta$, we know $x^{\prime}_{6}\sim_{\Delta}x_{3}$. Hence $x^{\prime}_{6}$ is common lower bound for $\{x_{1},x_{3},x_{5}\}$ in $(V\Delta^{\prime\prime},<)$.
Figure 13: Proof of Proposition 9.3
Now we assume there does not exists any vertex of type $\hat{b}_{1}$ such that it is adjacent in $\Delta$ to $\{x_{5},z,x_{1}\}$. In particular, $z_{1}\neq x_{6}$, $z_{5}\neq x_{6}$ and $z_{1}\neq z_{5}$. We claim there exists a vertex $x^{\prime}_{6}$ of type $\hat{b}_{3}$ such that it is adjacent in $\Delta$ to each of $\{z_{1},z_{5},x_{1},x_{5}\}$. Assuming the claim for the moment, then we apply Lemma 9.5 to the 8-cycles:
$\omega^{\prime\prime}=zz_{1}x_{1}m_{1}x_{2}m_{2}x_{3}z_{3},\ \omega^{\prime\prime\prime}=zz_{3}x_{3}m_{3}x_{4}m_{4}x_{5}z_{5}.$
See Figure 13 (II). If one of these two cycles are in Lemma 9.5 (1), then we are reduced to previous cases. Otherwise up to replacing $x_{2}$ and $x_{4}$ by different vertices of type $\hat{b}_{3}$, we can assume $x_{2}\sim_{\Delta}\{x_{1},z_{1},z_{3},x_{3}\}$ and $x_{4}\sim_{\Delta}\{x_{3},z_{3},z_{5},x_{5}\}$. Considering the $6$-cycle $x^{\prime}_{6}z_{1}x_{2}z_{3}x_{4}z_{5}$ in $\Delta$, we know $\{x^{\prime}_{6},x_{2},x_{4}\}$ are pairwisely upper bounded in $(V\Delta^{\prime},<)$, hence they have a common upper bound, denoted by $w$, in $(V\Delta^{\prime},<)$. We can assume $t(w)=4$. If $\{x^{\prime}_{6},x_{2},x_{4}\}$ are pairwise distinct, then by applying Lemma 9.4 to the 4-cycles $x^{\prime}_{6}x_{1}x_{2}w$, $wx_{2}x_{3}x_{4}$ and $wx_{4}x_{5}x^{\prime}_{6}$ in $\Delta$, we know $w$ is adjacent in $\Delta$ to $x_{i}$ for $i=1,3,5$. Thus $w$ is a common lower bound for $\{x_{1},x_{3},x_{5}\}$ in $(V\Delta^{\prime\prime},<)$. If at least two of $\{x^{\prime}_{6},x_{2},x_{4}\}$ is the same, say $x_{2}=x_{4}$, then $x_{2}$ is a common lower bound for $\{x_{1},x_{3},x_{5}\}$ in $(V\Delta^{\prime\prime},<)$.
It remains to prove the claim. See Figure 13 (III). As $\{x_{6},z_{1},z_{5}\}$ is pairwisely lower bounded in $(V\Delta^{\prime},<)$, Lemma 5.1 and Proposition 8.3 imply they have a common lower bound in the poset $(V\Delta^{\prime},<)$, denoted by $z_{6}$. Moreover, we can require $t(z_{6})=1$. As there does not exist any vertex of type $\hat{b}_{1}$ adjacent in $\Delta$ to each of $\{x_{5},x_{1},z\}$, we know $z_{6}\neq z$, $z_{6}\neq x_{1}$, $z_{6}\neq x_{5}$, $z_{1}\neq z_{5}$, $z_{1}\neq x_{6}$ and $z_{5}\neq x_{6}$. If $\operatorname{Type}(z_{6})=\hat{b}_{3}$, by applying Lemma 9.4 to the embedded 4-cycles $z_{6}x_{6}x_{1}z_{1}$, $z_{6}z_{1}zz_{5}$ and $z_{6}z_{5}x_{5}x_{6}$, we know $z_{6}$ is adjacent in $\Delta$ to each of $\{x_{1},x_{5}\}$. Thus the claim follows. Now assume $\operatorname{Type}(z_{6})=\hat{b}_{2}$. We will show this is impossible. By considering the 4-cycles $z_{6}x_{6}x_{1}z_{1}$ and $z_{6}z_{5}x_{5}x_{6}$, we deduce from Proposition 9.1 (1) that there is a vertex $w_{1}$ of type $\hat{a}$ adjacent in $\Delta$ to each vertex of $z_{6}x_{6}x_{1}z_{1}$, and a vertex $w_{5}$ of type $\hat{a}$ adjacent in $\Delta$ to each vertex of $z_{6}z_{5}x_{5}x_{6}$. Now $\{w_{1},w_{5},z\}$ is pairwisely upper bounded in $(V\Delta^{\prime},<)$, then they have a common upper bound, denoted $x^{\prime}_{6}$. As $w_{1}\neq w_{5}$ (otherwise $x_{1}\sim_{\Delta}z_{5}$, hence $z_{5}$ is adjacent to each of $x_{1},x_{5},z$, which was forbidden), we know $t(x^{\prime}_{6})=4$. Applying the bowtie free property of $(V\Delta^{\prime},<)$ to $\{x_{6},w_{1},x^{\prime}_{6},w_{5}\}$, we know $x_{6}=x^{\prime}_{6}$. As $x^{\prime}_{6}\sim_{\Delta}z$, we have $x_{6}\sim_{\Delta}z$. Hence $x_{6}$ is adjacent to each of $x_{1},x_{5},z$, which is forbidden.
Case 4: $\operatorname{Type}(x_{i})=\hat{b}_{3}$ for $i=2,4,6$. Now we consider the following three 8-cycles:
If one of these 8-cycles belong to Lemma 9.5 (1), then we are reduced to previous cases. If none of them belong to Lemma 9.5 (1), then for $i=2,4,6$, up to possibly replacing $x_{i}$ by another vertex of type $\hat{b}_{3}$, we can assume $x_{i}\sim_{\Delta}\{x_{i-1},z_{i-1},x_{i+1},z_{i+1}\}$. Now we repeat the argument in Case 3 to finish the proof. ∎
@@ 10 $K(\pi,1)$-conjecture for some 3D Artin groups
The goal of this section is to prove Corollary 10.4. We start with a preparatory lemma.
Lemma 10.1.
([Hua24a, Lemma 6.6]) Suppose $\Lambda^{\prime}$ is an admissible linear subgraph of the Coxeter diagram $\Lambda$ of $A_{S}$ and suppose the consecutive nodes of $\Lambda^{\prime}$ are $S^{\prime}=\{s_{i}\}_{i=1}^{n}$. Let $\Delta$ be the $(\Lambda,\Lambda^{\prime})$-relative Artin complex. Let $V$ be the vertex set of $\Delta$. We fix an order on $S^{\prime}$ such that it is compatible with one of the two linear orders on $\Lambda^{\prime}$, and endow $V$ with the induced order. Let $\Lambda^{\prime\prime}$ be a linear subgraph of $\Lambda^{\prime}$. Then the following holds.
  1. 1.
  2. 2.
    If $\Lambda^{\prime\prime}$ contains the largest (resp. smallest) node of $\Lambda^{\prime}$ and $\Delta_{\Lambda,\Lambda^{\prime}}$ is upward (resp. downward) flag, then $\Delta_{\Lambda,\Lambda^{\prime\prime}}$ is upward (resp. downward) flag.
  3. 3.
    Suppose $\Lambda^{\prime}$ only have two nodes. If $\Delta_{\Lambda,\Lambda^{\prime}}$ is bowtie free, then it is a graph with girth $\geq 6$. If in addition $\Delta_{\Lambda,\Lambda^{\prime}}$ is upward flag or downward flag, then it is a graph with girth $\geq 8$.
Proof.
@@
Let $A_{S}$ be an Artin group with $|S|=4$ which is free-of-infinity (i.e. its Coxeter diagram does not have $\infty$-labeled edges). If there exists $S^{\prime}\subset S$ with $|S^{\prime}|=3$ such that $A_{S^{\prime}}$ is not spherical, then the Artin complex for $A_{S}$ is contractible.
@@
Let $S=\{a,b,c,d\}$ and $S^{\prime}=\{a,b,c\}$. Let $\Delta_{S}$ (resp. $\Delta_{S^{\prime}}$) be the Artin complex for $A_{S}$ (resp. $A_{S^{\prime}}$). By Lemma 2.2, the link of each vertex of type $\hat{d}$ in $\Delta_{S}$ is isomorphic to $\Delta_{S^{\prime}}$. As $A_{S^{\prime}}$ is not spherical, it is a 2-dimensional Artin group. It is known the modified Deligne complex, defined in [CD95], for any 2-dimensional Artin group is contractible [CD95]; and the modified Deligne complex isomorphic to the barycentric subdivision of $\Delta_{S^{\prime}}$. Thus the link of each vertex of type $\hat{d}$ is contractible. Then $\Delta_{S}$ is homotopic equivalent to the relative Artin complex $\Delta^{\prime}=\Delta_{S,S^{\prime}}$ by Lemma 2.4.
@@
Now we look at the case $A_{bcd}$ is not spherical. Then $\operatorname{lk}(w,\Delta_{S})$ quotiented by the action of $A_{bcd}$ is a triangle $T^{\prime}$, whose vertices are denoted by $\hat{b},\hat{c},\hat{d}$. As $\frac{1}{m_{bc}}+\frac{1}{m_{cd}}+\frac{1}{m_{bd}}\leq 1$, we can realize $T^{\prime}$ as a geodesic triangle in $\mathbb{E}^{2}$ or $\mathbb{H}^{2}$ such that the angle at $\hat{b}$ is $\frac{\pi}{m_{cd}}$, the angle at $\hat{c}$ is $\frac{\pi}{m_{bd}}$ and the angle at $\hat{d}$ is $\frac{\pi}{m_{bc}}$. This induces a metric on $\operatorname{lk}(w,\Delta_{S})$ which is known to be CAT$(0)$ (cf. [CD95]). As $\operatorname{lk}(w,\Delta^{\prime})$ is the full subcomplex of $\operatorname{lk}(w,\Delta)$ spanned by vertices of type $\hat{b}$ and $\hat{c}$, we know $\operatorname{lk}(w,\Delta^{\prime})$ has girth $2m_{bc}$ by [Hua24a, Lemma 9.8 (1)]. Thus the claim is proved. ∎
@@
Suppose $A_{S}$ is an Artin group with $|S|\leq 4$. Assume that its Coxeter diagram is not a $(3,5,3)$-linear diagram. Then $A_{S}$ satisfies the $K(\pi,1)$-conjecture.
@@
We assume $A_{S}$ is irreducible. By [ES10], it suffices to consider the case when $A_{S}$ is free-of-infinity. We also assume $A_{S}$ is not spherical, otherwise the result follows from [Del72]. When $|S|\leq 3$, $A_{S}$ is 2-dimensional and the theorem follows from [CD95]. Now we assume $|S|=4$. If there exists a subset $S^{\prime}\subset S$ such that $|S^{\prime}|=3$ and $A_{S^{\prime}}$ is not spherical, then the theorem follows from Proposition 10.2. It remains to consider the case when $|S|=4$ and $A_{S^{\prime}}$ is spherical whenever $S^{\prime}$ is a 3-element subset of $S$. By a result of Lanner [Lan50], $A_{S}$ is either an affine Artin group, or a 3-dimensional hyperbolic cocompact tetrahedron group, i.e. its associated Coxeter group acts on $\mathbb{H}^{3}$ properly and cocompactly by isometries such that its fundamental domain is a tetrahedron. The $K(\pi,1)$-conjecture for affine Artin groups are proved in [PS21]. For the 3-dimensional hyperbolic tetrahedron groups, there are only nine of them, five has Coxeter diagram being a cycle, which is treated in [Garside, Hua24a]. Then remaining four groups have Coxeter diagrams as in Figure 14. By Theorem 2.5, it suffices to show the associated Artin complex is contractible.
@@
By Theorem 2.5, it suffices to show the Artin complexes associated with these Coxeter diagrams are contractible. First we look at the case of $(5,3,4)$ and $(5,3,5)$ linear diagrams. Assume the nodes in the Coxeter diagram are ordered from left to right, and we endow the vertex set of the associated Artin complex $\Delta$ with the induced order. Then the contractibility of $\Delta$ follows from Theorem 2.18, Theorem 7.1, Theorem 2.37, Theorem 2.35 and Lemma 10.1. In the case when $\Lambda$ is a tripod, we deduce the contractibility of the Artin complex from Proposition 8.3, Theorem 7.1, Theorem 5.6, and Theorem 2.35. ∎
@@
Let $W_{S}$ be a reflection group acting properly on $\mathbb{H}^{3}$ with finite volume fundamental domain. Assume the Coxeter diagram $\Lambda$ of $W_{S}$ is not the linear graph with consecutive edges labeled by $(3,5,3)$. Then the $K(\pi,1)$-conjecture holds true for the associated Artin group $A_{S}$.